OpenStudy (firejay5):

Simplify. 23. 3 cubed root of 56y^6z^3 24. 2 cubed root of 24m^4n5 25. 4th root of 1/81c^5d^4 26. 5th root of 1/32w^6z^7

OpenStudy (anonymous):

I'm sorry don't really know this? What lesson is it? Maybe this can help: http://www.wolframalpha.com/input/?i=3+cuberoot+of+56y%5e6z%5e3&incTime=true

OpenStudy (firejay5):

23. 3\[\sqrt[3]{56y^6z^3}\] 24. 2 \[\sqrt[3]{24m^4n^5}\] 25.\[\sqrt[4]{\frac{ 1 }{ 81 }} c^5d^4\] 26. \[\sqrt[5]{\frac{ 1 }{ 32 }}w^6z^7\]

OpenStudy (firejay5):

I meant on 23 and 24 it's cubed and the number 23. 3 goes with the whole radical expression and 24 the 2 goes with its whole radical expression

OpenStudy (campbell_st):

here is some advice for the 1st problem\[\sqrt[3]{7} \times (2^3)^{\frac{1}{3}}\times(y^6)^{\frac{1}{3}} \times(x^3)^{\frac{1}{3}}\] use the index la to simplify

OpenStudy (firejay5):

I am on #26 now

OpenStudy (firejay5):

@campbell_st How would you write #25 and 26 on http://www.wolframalpha.com/

OpenStudy (firejay5):

Do you know @Hero

hero (hero):

That isn't the right question to ask. Of course I know. The question is will I drop what I'm doing right now to help.

OpenStudy (firejay5):

Just real quick; could you enter them in on the website and copy and paste the question of it for both problems